Need answer confirmation please.
Q: Write equation of line through (2,-3) parallel to 2x-3y=0. Write equation is standard form.
What I did:
2x-3y=0 <=== I found a slope of 2/3
Then:
y-(-3)=2/3(x-2)
y+3=2/3(x-2)
Then to Standard form:
y+3=2/3x-4
y+3-3=2/3x-4-3
y=2/3x-7
From there:
I subtracted -2/3 from both sides and arrived at: -2/3x+y=-7
Is this correct?
Thank you!

Answer by solver91311(24713)



Your answer in italics as far as you were correct:

2x-3y=0 <=== I found a slope of 2/3

Then:

y-(-3)=2/3(x-2)

y+3=2/3(x-2)


Then you started to go awry:

Your next step should have been to get rid of the fractional slope by multiplying by the denominator, namely 3:





Distribute:



Put the variable terms in the LHS and the constants in the RHS



Collect



I like my lead coefficient to be positive, so:




John

My calculator said it, I believe it, that settles it